OpenStudy (anonymous):

Question A canoe can go 24 km downstream in 3 hours. The return trip takes 4 hours. What is the speed of the current?

OpenStudy (anonymous):

boat speed =a current=c 3(a+c)=24 4(a-c)=24

OpenStudy (anonymous):

3a+3c=24 4a-4c=24 solve for a and c

OpenStudy (anonymous):

so i am supposed to add these two equations together im lost??

OpenStudy (anonymous):

How would you solve system of equations

OpenStudy (anonymous):

PEMDAS- would i use that??

OpenStudy (anonymous):

Easy jst divide by 3 both sidez 3(a+c)=24 yu will get a+c=8 Then divide by 4 both sidez 4(a-c)=24 yu will get a-c=6 a+c=8 a-c=6 then yu will get 2a=14 then a=7 c=1

OpenStudy (anonymous):

Here is a step by step solution: Let Vd = speed downstream Vu = speed upstream Vc = speed of the canoe Vs = speed of the stream Going downstream we know Vd = 24/3 = 8. Going upstream we know Vu = 24/4 = 6. Going downstream we are going in the direction of Vs so: Vd = Vs + Vc or 8=Vs + Vc Going upstream we are going against the direction of Vs so: Vu =Vc - Vs or 6 = Vc - Vs So we have two equations in two unknowns. Solve them: Vc+Vs = 8 Vc-Vs = 6 ----------- 2Vc = 14 Vc = 7 Vs = 1